OpenStudy (anonymous):

the electric potential in a region of space is \[V=350/\sqrt{x ^{2}+y ^{2}}\] where x and y are in meters. what is the strength of electric field at (x,y)=(2.6,2.8)m?

OpenStudy (anonymous):

E=-dV/ds. gradient V= 350(-1/2)(x^2+y^2)^(-3/2)(2x)+350(-1/2)(x^2+y^2)^(-3/2)(2y). anything wrong so far?

OpenStudy (anonymous):

That kinda how i did it but I keep getting the wrong answer. Let me know if this formula worked for you

OpenStudy (anonymous):

my answer's not working either =/

OpenStudy (turingtest):

wrong so far is that you have written the gradient as a scalar when it is a vector

OpenStudy (turingtest):

\[\vec E=-\nabla V=<-350x(x^2+y^2)^{-3/2},-350y(x^2+y^2)^{-3/2}>\]now plug in the variables and find the magnitude of the vector

OpenStudy (turingtest):

the magnitude is given by\[\large |E|=\sqrt{E_x^2+E_y^2}\]

OpenStudy (anonymous):

thank you

OpenStudy (turingtest):

welcome, I hope it works

OpenStudy (anonymous):

worked!!

OpenStudy (turingtest):

sweet :D

OpenStudy (anonymous):

If you cartesian coordiante to spherical coordiante, you will find the answer easier.

OpenStudy (anonymous):

o yea but how? use r=sqrt(x^2+y^2) and then......? can't believe i can't remember how